### v1

Abstract
We demonstrate, for the first time, that neural scaling laws can deliver near-exact solutions to the many-electron Schrödinger equation across a broad range of realistic molecules. This progress is enabled by the Lookahead Variational Algorithm (LAVA), an effective optimization scheme that systematically translates increased model size and computational resources into greatly improved energy accuracy for neural network wavefunctions. Across all tested cases, including benzene, the absolute energy error exhibits a systematic power-law decay with respect to model capacity and computation resources. The resulting energies not only surpass the 1 kcal/mol “chemical-accuracy” threshold but also achieve 1 kJ/mol subchemical accuracy. Beyond energies, the scaled-up neural network also yields better wavefunctions with improved physical symmetries, alongside accurate electron densities, dipole moments, and other important properties. Our approach offers a promising way forward to addressing many long-standing challenges in quantum chemistry. For instance, we improve energetic properties for systems such as the potential energy curve of nitrogen dimer as dissociation is approached and the cyclobutadiene automerization reaction barrier, producing definitive benchmarks, particularly in regimes where experimental data are sparse or highly uncertain. We also shed light on the decades-old puzzle of the cyclic ozone stability with highly accurate calculations for the cyclic-to-open ozone barrier. These results provide near-exact reference calculations with unprecedented accuracy, universal reliability and practical applicability, establishing a foundation for AI-driven quantum chemistry.
1 Main
The many-electron Schrödinger equation lies at the foundation of quantum chemistry and condensed matter physics, providing a first-principles framework for understanding the quantum nature of the physical world. Despite its central importance, no general-purpose method has come close to solving it exactly for realistic systems. Instead, practical quantum chemistry has long relied on the cancellation of large and often uncontrolled errors to reach the so-called “chemical accuracy”–typically defined as energy errors within 1 kcal/mol for relevant energy differences. This scheme, however, comes with several significant limitations, including poor performance in predicting observables beyond relative energies and a lack of systematic error control. For instance, the widely used Density Functional Theory (DFT) can predict qualitatively incorrect electron densities [1], undermining its reliability for density-derived properties, such as dipole moments and polarizabilities. Correlated wavefunction methods, on the other hand, depend on error cancellation due to steep computational scaling and slow basis-set convergence [2, 3], compromising their reliability in complex problems. These limitations underscore the need for a more accurate and reliable solution that delivers cancellation-free, high-accuracy energies and other observables for realistic systems.
Neural network-based quantum Monte Carlo (NNQMC) has emerged as a promising ab initio wavefunction theory to this challenge [4, 5, 6]. Unlike other machine learning approaches in quantum chemistry that rely on precomputed labeled data (e.g., DFT energies) [7, 8], NNQMC obtains the target quantum state directly through unsupervised optimization, without requiring any reference data. In particular, the full many-body wavefunction is modeled with highly expressive neural networks, providing access to accurate total energies, high-quality wavefunctions, and derived observables including electron and spin densities. Recent progress has improved energy accuracy as well as computational efficiency, highlighting its potential as a next-generation quantum chemistry framework [9, 10, 11, 12, 13, 14]. Nonetheless, NNQMC has not yet come meaningfully close to the exact solutions of the Schrödinger equation as molecular systems increase in size and complexity. This is partly because default-sized neural networks lack sufficient representational capacity for larger systems, while simply increasing the network size rarely leads to proportional improvements in accuracy due to optimization challenges.
In this work, we train neural‑network wavefunctions that, for the first time, deliver near‑exact solutions to the many‑electron Schrödinger equation for realistic systems with up to 12 atoms, achieving accuracy on par with experimental uncertainty. The resulting absolute energies surpass traditional “chemical accuracy” to approach the 1 kJ/mol regime, thereby enabling definitive relative energies without relying on error cancellation. Additionally, these solutions provide accurate many-body wavefunctions, which in turn produce benchmark-quality physical observables including electron density and dipole moments. This capability is made possible by two key innovations. First, we introduce the Lookahead Variational Algorithm (LAVA), an optimization scheme that combines variational and projective frameworks, offering better performance over existing schemes in NNQMC, such as Variational Monte Carlo (VMC) with stochastic reconfiguration [15] and Wasserstein Quantum Monte Carlo (WQMC) [16]. Second, we present the first systematic study of neural scaling laws [17] in ab initio quantum chemistry, showing that total energy errors decrease systematically and predictably by simply scaling up neural network model capacity and computational resources. Together, these advances position our approach uniquely within quantum chemistry, offering near-exact total energies and wavefunctions at the full configuration interaction (FCI) accuracy and complete basis set (CBS) limit at the same time.
Through this approach, we address several long-standing challenges in quantum chemistry, demonstrating its universal accuracy, flexibility, and ability to provide definitive benchmarks. Firstly, we establish a high-quality benchmark for cyclobutadiene’s transition barrier, aligned with refined experimental data as well as the best estimates from coupled cluster (CC) and configuration interaction (CI) methods. Secondly, we study the potential energy curve (PEC) of diatomic molecules, which governs the vibrational levels critical for astrophysical models of planetary and stellar atmospheres [18]. In particular, we focus on nitrogen dimer and present a new PEC benchmark that surpasses previous experiment-based references in both accuracy and reliability. Thirdly, we reassess the metastability of the cyclic ozone, helping to provide a definitive answer to a long-standing controversy [19]. Collectively, these findings confirm our approach as a new standard of accuracy for predictive quantum chemistry.
2 Results
2.1 Scale Toward Exact Solutions with LAVA
Figure 1: Scaling Up LAVA-Optimized Neural Network Wavefunctions Toward Exactness. a. Upper: The many-electron wavefunction is modeled by a massively parameterized neural network and optimized with LAVA or neural network-based variational Monte Carlo (NNVMC). Middle: A conceptual energy landscape illustrates LAVA’s better convergence behavior throughout the training process, compared to NNVMC, which only considers energy-based losses. Lower: LAVA combines both the gradient of an energy-based loss and a projection-derived direction via the Lookahead algorithm. b. Main panel: By scaling up the network size and computation resources, LAVA achieves systematic and fool-proof improvements in accuracy, surpassing chemical accuracy and approaching exact solutions. In contrast, traditional computational methods, such as coupled cluster and configuration interaction, suffer from steep computational scaling and resource bottlenecks, due to various issues such as basis set limitations, exploded determinant growth, and steep scaling with higher excitations. Inset: A schematic plot of different quantum chemistry methods in terms of accuracy and system size. Notably, LAVA is able to achieve more accurate energy results than previous NNVMC works with the same neural network size.
Figure 2: Neural scaling laws enable breakthrough performance. a. LAVA delivers near-exact ground state energy for benzene as neural networks scale up. The neural scaling law here is described by $E-E_{\mathrm{expt}}\propto N_{p}^{-0.52}$, where $E$ denotes LAVA energies, $E_{\mathrm{expt}}$ denotes experiment-derived reference [20], $N_{p}$ is the number of parameters in the neural networks. We also include energies from NNVMC [9] and CCSD(T)/CBS [21] for comparison. b. Upper: Absolute energy errors of LAVA-SE, CCSDTQ5/CBS level W4-theory [22], CCSD(T)/CBS level W2-theory [22], and NNVMC. LAVA-SE is more accurate than 1 kJ/mol thresholds (purple shade), aligning closely with highly accurate W4 theory. Experimental references are based on atomization energies from ATcT (Active Thermochemical Tables) [22] and absolute atom energies from Chakravorty et al. [23]. Specifically, for atom F, we use our variationally lower result since it is more reliable. Details are in Supplementary Note 4.1. NNVMC uses LapNet [9] with the default model size. Lower: Neural scaling laws of LAVA on the same molecules, showing error reduction with model size in power-law trends. See Supplementary Note 2 for details. c. Cost scaling exponents for N2 as LAVA increases model size and CC methods increase excitation orders. For CC, data from Chan et al. [24] shows that CC’s computational effort scales as a power-law trend relative to $1/(E_{\mathrm{CC}}-E_{\mathrm{FCI}})$ under cc-pVDZ basis set. For LAVA, the total GPU time cost scales as a power-law relative to $1/(E_{\mathrm{LAVA}}-E_{\mathrm{SE}})$, where $E_{\mathrm{SE}}$ denotes the LAVA-SE result. LAVA maintains a nearly constant exponent along the $N_{2}$ dissociation curve, while CC’s performance deteriorates in the strongly correlated region. d. LAVA GPU hours as the number of electrons $N_{e}$ increases, with increasing model size for better accuracy. With model size fixed, runtime scales as $N_{e}^{2.1}$ (dotted lines). To ensure chemical accuracy, runtime scales as $N_{e}^{5.3}$ (the solid line).
In this section, we present a systematic study of neural scaling laws that achieve beyond-chemical-accuracy solutions to the Schrödinger equation. We demonstrate that neural network-based solutions systematically approach exact results with increasing model capacity and training compute. Notably, the convergence curve of absolute energy error follows a robust power-law decay, and the same convergence behavior is consistently observed across diverse molecular systems.
Although scaling laws have reshaped various domains of Artificial Intelligence [25, 26, 27, 28, 29, 30, 31, 32], their application within quantum science remains underexplored [33]. More critically, the benefit of scaling up neural‑network wavefunctions appears limited: energy improvement tends to saturate well before reaching exactness, hinting at an inability to fully exploit the capacity of large neural networks. To overcome this bottleneck, we introduce LAVA, an improved optimization framework for neural network wavefunctions that combines variational Monte Carlo updates with a projective step inspired by imaginary time evolution (Fig. 1a). This two-step procedure is effective for eluding local minima during the neural network training process, which is crucial for achieving asymptotic exactness as the neural network ansatz scales up. In practice, LAVA significantly improves stability during the training process and accuracy at the end of training, resulting in better wavefunctions and energies. See Section 4.2 and Supplementary Note 1.1 for algorithmic details.
LAVA scales predictably and robustly to deliver accurate solutions beyond chemical accuracy. In pursuing systematic improvement towards exactness, our approach offers several key advantages. Most notably, LAVA requires little heuristic tuning or chemical intuition, making it effectively a “fool-proof” process. Moreover, LAVA avoids the prohibitive scaling with excitation order inherent to traditional methods such as coupled cluster theory, offering a significantly more efficient route to high accuracy. These advantages are sketched in the main panel of Fig. 1b.
To demonstrate the effectiveness of neural scaling laws, we provide a range of quantitative evidence, as illustrated in Fig. 2. We examine the absolute energy of representative organic molecules, for which highly accurate experimental benchmarks (via total atomization energy) and theoretical references (W4 theory) are available. Our energy estimates not only surpass chemical accuracy but also fall within experimental uncertainty, a more stringent criterion for accuracy. Fig. 2a illustrates how neural scaling laws enable such accuracy, using benzene as a representative example. Specifically, the energy error decays following a power-law relation as the number of parameters in the neural network increases. In addition, we observe a linear relation between energy and variance as neural networks scale up (See Supplementary Note 3). Accordingly, we adopt an energy-variance extrapolation scheme (Section 4.4), which yields our best energy estimate, LAVA with Scaling-law Extrapolation (LAVA-SE).
In all tested cases, we consistently achieve total energy accuracy at sub-kJ/mol level using LAVA-SE together with a robust and reproducible scaling pattern of LAVA (Fig. 2b). Remarkably, our results align closely with those of the W4 protocol, a CC-based composite procedure extrapolated to the all-electron CCSDTQ5/CBS energy, offering kJ/mol or even “semi-spectroscopic” accuracy for thermochemistry [22]. With LAVA, applying neural scaling laws alone suffices to surpass the sub-kJ/mol threshold, without the need for ad hoc corrections. Note that this level of accuracy is achieved without relying on any error cancellation, offering a direct and absolute measure of proximity to the exact solution of the Schrödinger equation.
We further assess the efficiency and practicality of our approach from two complementary perspectives, related to other scaling behaviors that emerge as we scale up our calculations. First we demonstrate that LAVA can maintain favorable convergence speed across both strongly or weakly correlated regimes, taking nitrogen molecule dissociation curve as a representative case in Fig. 2c. Specifically, in order to approach the exact solution (using the LAVA-SE as a reference), the computation scaling with respect to computational runtime remains close to linear with minor fluctuations along the whole curve (Fig. 2c lower panel). As a comparison, couple cluster (CC) theory exhibit significant deterioration in convergence scaling in the strongly correlated regime [24] (Fig. 2c upper panel). Moreover, LAVA’s scalability can further benefit from parallel computing, whereas CC cannot (see Supplementary Note 2.2). Second, we examine LAVA’s performance as the molecular size increases. These results demonstrate that our approach is able to maintain chemical accuracy with a relatively low computational scaling ($N_{e}^{5.2}$, where $N_{e}$ denotes the number of electrons), as shown in Fig. 2d. By comparison, CCSD(T) scales as $N_{e}^{7}$, while achieving chemical accuracy may require even higher-order excitations, which impose steeper, if not impractical, scalings.
In the following sections, we address a range of practical chemical challenges using scaled‑up neural‑network wavefunctions. When a variational guarantee is preferred, we report energies from our largest network; otherwise, we use the extrapolated LAVA‑SE results.
2.2 Definitive Benchmarks beyond Experimental Limitations
Figure 3: Definitive LAVA benchmarks complement experiments. a. Upper: Reaction barriers of $\mathrm{C}_{4}\mathrm{H}_{4}$ derived from experiment and different theoretical methods. Bars represent the spread of reported values, corresponding to each line, within each method family. Stars indicate the improved experimental estimate and the best estimates from systematically improvable theories, namely CCSDTQ/aug-cc-pVTZ in CC, HCI/cc-pVTZ extrapolated to the FCI limit in selected CI, and LAVA in NNQMC. The dashed line shows the value predicted by LAVA, highlighting the consistency among those stars. Lower: Power-law scaling trends between the number of parameters $N_{p}$ and energy difference $E-E_{\mathrm{SE}}$, where $E_{\mathrm{SE}}$ are scaling-law extrapolation results. b. Upper main panel: The ground-state potential energy curve of N2 dissociation. The analytic potential functions, namely $\text{MLR}_{3}(9)$, $\text{MLR}_{4}(6,8)$, and $\text{EMO}_{2}(6)$, are fitted from vibrational levels $v=0-19$. Upper inset: PECs fitted from different analytic function forms and parameters show a large discrepancy at long bond lengths. Lower main panel: The energy difference between two ab initio calculations and three fitted analytic PECs. $\text{MLR}_{3}(9)$, the surrogate model of LAVA, is the new recommended benchmark. The $r_{12}$-MR-ACPF result is shifted together so that its results match high-fidelity experimental data in the equilibrium region (the shaded area).
The scarcity of high-quality experimental data remains a key bottleneck in quantum chemistry, limiting both the development of exchange-correlation (XC) functionals in DFT and the validation of advanced wavefunction methods. For unstable or non-equilibrium geometries, it is challenging to perform thermochemical experiments to measure their energies and observables. In such cases, it is particularly valuable to have highly accurate ab initio methods to fill the gaps. LAVA is well-suited for these challenging regimes, offering a reliable and systematically improvable alternative that serves as a critical complement to experimental data.
As a first demonstration, we study the reaction barrier associated with the automerization of cyclobutadiene ($\mathrm{C}_{4}\mathrm{H}_{4}$), a long-standing challenge in quantum chemistry due to its multireference character. To date, neither experimentally derived benchmarks nor theoretical predictions can reliably determine the reaction barrier between the rectangular ($D_{\text{2h}}$) minimum and the square ($D_{\text{4h}}$) transition state. The experimental estimate gave a lower bound of 1.6 kcal/mol and an upper bound of 10 kcal/mol, while various theoretical predictions range from 3 to 20 kcal/mol [34, 35, 36, 37, 5, 21], as illustrated in Fig. 3a. Here, we establish a definitive benchmark for this transition barrier (9.2 kcal/mol), obtained from LAVA-SE for both $D_{\text{2h}}$ and $D_{\text{4h}}$ states. This result is derived from a sequence of scaled-up neural network wavefunctions and does not rely on error cancellation. To further validate our result, we establish consensus with improved experimental estimates and the most accurate predictions from different theoretical frameworks, as shown by the star symbols in Fig. 3a. For experimental data, we refined the previous rough estimate to a more precise value of 9.9 kcal/mol (See Supplementary Note 6). This new estimate agrees well with LAVA and may align even more closely under alternative computational settings during refinement [38]. On the theoretical side, we leverage the systematic improvability of the CC and selected CI families to obtain their best estimates, which are CCSDTQ/AVTZ [34] and HCI(20e, 172o)/cc-pVTZ extrapolated to the FCI limit [36], respectively. Notably, the best estimates from NNQMC, CC, and selected CI—three fundamentally different theoretical approaches—agree with each other within 0.3 kcal/mol, suggesting convergence to the exact reaction barrier. This consistency highlights LAVA’s ability to deliver accurate and reliable solutions for challenging electronic structure problems.
Next, we demonstrate LAVA on the potential energy curve of the nitrogen dimer, a prototypical and long-standing challenge due to its multireference character, and establish a new benchmark, MLR3(9), that surpasses the previous standard MLR4(6,8) in accuracy. The previous benchmark, labeled as MLR4(6,8) in Fig. 3.b, was fitted to experimental vibrational levels $v=0-19$ [39]. These vibrational levels provide a high-accuracy energy benchmark only in the near-equilibrium region, corresponding to bond lengths $r$ between 0.9 to 1.5 $\mathrm{\AA }$ (gray shaded region). There are few experimental results for vibrational levels $v$=20-25 [40], albeit with a high uncertainty. Measuring higher vibrational levels is challenging because the nitrogen dimer becomes unstable in the near-dissociation region.
Different fitting schemes also give large discrepancies in this region. As shown in the inset of Fig. 3, EMO2(6) and MLR4(6,8) differ by 3.4 kcal/mol in the near-dissociation region, where experimental benchmarks are unavailable. Details of fitting analytic PECS from experimental vibrational levels are given in Supplementary Note 7 and Le Roy et al. [39]. At the intermediate region, MLR4(6,8) also shows significant disagreements with multi-reference correlated calculations [41] and previous NNQMC studies [6, 42, 43, 44, 21].
[…]
Collectively, the cases in this section demonstrate the reliability and versatility of our approach. It not only resolves long-standing discrepancies between experimental measurements and theoretical predictions but also provides definitive benchmarks when experiments or conventional theories fall short, thereby establishing a new standard for future benchmark studies.
2.3 Metastability of Ozone’s Ring-Minimum Species
Figure 4: Analysis of ozone bent-cyclic isomerization reaction. a. A conceptual diagram for $\mathrm{O}_{3}$ potential energy surfaces of ${}^{1}\mathrm{A}^{\prime}$, ${}^{1}\mathrm{A}^{\prime\prime}$, ${}^{3}\mathrm{A}^{\prime\prime}$, and $2^{1}\mathrm{A}^{\prime}$ in isosceles triangle geometries, based on XMS-CASPT2 predictions [46, 47]. b. Summary of calculated energy barriers of $\mathrm{O}_{3}$ isomerization reaction from ring-minimum species to open-ring minimum since 1972. Our LAVA-SE estimate, denoted by the purple star, is consistent with most high-accuracy results in the last decade. c. Neural scaling laws for OM, TS, RM configurations of ozone, respectively, showing clear power-law decay between energy error and network size. The energy error is calculated with respect to LAVA-SE. d. Evolution of wavefunction spatial symmetry as the number of parameters increases for the TS configuration. $\sigma_{h}$ is the expectation of the horizontal reflection operator. With the default network size, LAVA produces a contaminated state as a superposition of 1A state (weight 0.18) and 1A′′ state (weight 0.81). Enlarged networks yield a pure ${}^{1}\mathrm{A}^{\prime\prime}$ state, which is 3 kcal/mol lower in energy compared to the contaminated one. For validation purpose, we also generate symmetry-enforced LAVA results for 1A and 1A′′, visualizing their natural orbitals with the lowest two occupation numbers that $\mathrm{occ}>0.1$.
Ozone plays a critical role in atmospheric chemistry. A long-standing puzzle, however, is the metastability of cyclic ozone, which has been predicted by various theoretical studies, from early ab initio work to modern coupled-cluster and multireference methods, since the 1970s [48, 49, 19, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. Nonetheless, direct experimental evidence remains absent, casting uncertainty on these theoretical predictions.
[…]
Additional evidence further supports the accuracy and reliability of our results. To begin with, the predictions of LAVA regarding the energy and geometry of OM are in excellent agreement with experimental results [61] (See Supplementary Note 8.1). Furthermore, neural scaling laws enable the emergence of physical symmetries without applying explicit constraints to wavefunctions. As the model size increases, the neural network wavefunction recovers the correct spin multiplicity and spatial symmetry. Take TS as a showcase, LAVA guides the neural network wavefunctions toward the correct ${}^{3}\mathrm{A}^{\prime\prime}$ ground state without any constraints (see Supplementary Note 8.6). Additionally, when we enforce singlet spin symmetry using a penalty-based method [10], LAVA identifies the ${}^{1}\mathrm{A}^{\prime\prime}$ ground state, as evidenced by the improved spatial symmetry depicted in Fig. 4d.
Overall, such a capability of scaling up towards the correct ground state is essential for reliable ground-state characterization in strongly correlated regimes, where multiple near-degenerate states often engage in competition.
2.4 Beyond Energies
Figure 5: LAVA delivers reliable dipole moments. a. Dipole moments $\mu$ derived from our wavefunctions match experimental values $\mu_{\mathrm{expt}}$ [62]. For comparison, CCSD(T)/CBS results from Hait and Head-Gordon [63] are also visualized, which fall out of the experimental range for NO, O3, and F2O2. For all molecules, the SCAN density functional results [63] fail to match experimental references. b. Basis set convergence of dipole moment of F2O2 for HF and coupled-cluster methods, together with experimental benchmark and LAVA estimates for comparison. Dipole moments in Debyes are plotted against correlated-consistent basis sets (cc-pVXZ) and augmented counterparts (aug-cc-pVXZ) for $\mathrm{X}=\mathrm{D},\mathrm{T},\mathrm{Q},5$. CBS extrapolation scheme follows Supplementary Note 6.2. The missing points indicate calculations infeasible with available computational resources. Results of coupled-cluster methods exhibit strong basis set dependence and severe memory bottlenecks, while LAVA achieves experimental-level accuracy.
Beyond accurate absolute energies, LAVA also produces near-exact wavefunctions in the CBS limit, benefiting from its first-quantized nature. Consequently, it delivers FCI/CBS-quality physical observables including electron densities and dipole moments. Conventional quantum chemistry methods, by contrast, remain limited by the finite basis set approximation, despite decades of effort toward developing systematically improvable basis sets [64]. While relative energies tend to converge rapidly with basis set size [65, 66], the convergence behavior of electron densities and density-derived properties toward the CBS limit remains less well understood [63, 67].
We assess LAVA’s performance on dipole moments, a critical property reflecting molecular charge distribution and polarity. Specifically, we compare LAVA predictions with highly accurate references from spectroscopy experiments [62] (Fig. 5a). Across all the molecules examined, LAVA shows excellent agreement with experimental data, falling within the bounds of experimental uncertainty. In contrast, CCSD(T)/CBS results deviate from these bounds for molecules with significant multireference character (see Supplementary Table 17), namely $\text{O}_{3}$ and $\text{F}_{2}\text{O}_{2}$. One could, in principle, move beyond CCSD(T)/CBS to CCSDT(Q)/CBS for better accuracy, but such calculations are computationally intractable due to the steep scaling of coupled cluster methods. Moreover, dipole moments are notoriously sensitive to the choice of basis set and converge slowly toward the CBS limit. Fig. 5b exemplifies this behavior for $\text{F}_{2}\text{O}_{2}$, where neither CCSD nor CCSD(T) reaches experimental accuracy when extrapolated along either the correlated-consistent basis set sequence (cc-pVXZ) or its augmented counterpart (aug-cc-pVXZ). Memory demand grows rapidly with increasing basis set size, emerging as the computational bottleneck: CPU-based PySCF [68] is limited to the aVTZ basis set, while GPU-accelerated ByteQC [69] can accommodate the larger aV5Z basis set. Notably, LAVA also produces more accurate energy than CC-based composite theory W4 for molecules with strong multireference character, as discussed Supplementary Note 4.3.
Moreover, with scaled-up neural networks, LAVA is also able to produce high-quality electron densities [70]. Together with benchmark-quality energies, this provides critical reference data for developing better density-functional approximations that seek to recover the exact energy from the exact density [1].
3 Discussion
This study demonstrates how LAVA-powered neural scaling laws can systematically and practically approach exact solutions to many-electron Schrödinger equations. We generate high-quality benchmarks for various molecules, overcoming the issues of uncertainties and inaccuracies in experimental data. Additionally, we confirm the metastability of ring-minimum ozone from a theoretical perspective.
We are now able to provide FCI/CBS quality wavefunctions with an attractive computational scaling. This enables the creation of a reliable benchmark dataset of both energies and densities, complementing scarce experimental references. Such datasets can facilitate data-driven development of next-generation XC functionals in DFT and rigorous validation of other state-of-the-art wavefunction methods, opening new chapters in high-accuracy quantum chemistry and its applications. Besides its favorable computational scaling and embarrassingly parallelism, LAVA also benefits from other recent algorithmic progresses in the field of NNQMC [12, 71, 13], which further reduce the computational cost significantly. Moreover, with rapid advances in AI hardware and optimization techniques, we anticipate that this approach will become increasingly practical and scalable, enabling broader applications to larger systems in quantum chemistry and beyond.
Overall, the synergy of AI and quantum chemistry offers a powerful route to solving complex electronic structure problems with near-exact solutions, unlocking transformative applications in catalysis, materials science, and drug discovery. While challenges remain, neural scaling laws establish a robust foundation for advancing accuracy and scalability across increasingly complex systems in ab initio quantum chemistry.
4 Methods
4.1 Variational Optimization Framework
In this work, we focus on the time-independent electronic Schrödinger equation within the Born-Oppenheimer approximation:
$\hat{H}\psi$ $\left({x}_{1},\ldots,{x}_{N}\right)=E\psi\left({x}_{1},\ldots,{x}_{N}\right),$ (1)
$\hat{H}=$ $-\frac{1}{2}\sum_{i}\Delta_{i}+\sum_{i>j}\frac{1}{\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|}$
$-\sum_{iI}\frac{Z_{I}}{\left|\mathbf{r}_{i}-\mathbf{R}_{I}\right|}+\sum_{I>J}\frac{Z_{I}Z_{J}}{\left|\mathbf{R}_{I}-\mathbf{R}_{J}\right|},$ (2)
where for the $i$-th electron ($i\in\{1,2,\cdots,N\}$), ${x}_{i}=\{\mathbf{r}_{i},\sigma_{i}\}$ consists of the coordinate of electron $\mathbf{r}_{i}\in\mathbb{R}^{3}$ and its spin $\sigma_{i}\in\{1,-1\}$, and for the $I$-th nucleus ($I\in\{1,2,\cdots,M\}$), we denote the charge as $Z_{I}$ and its fixed position as $\mathbf{R}_{I}$. Obeying the spin-statistics theorem, a many-electron wavefunction is antisymmetric under an exchange of electrons.
As an ab initio method, variational Monte Carlo (VMC) directly solves the following optimization problem to approximate the ground state of a many-electron quantum system:
$\mathop{\mathrm{min}}\limits_{\theta}$ $\qquad\frac{\langle\psi_{\theta}|\hat{H}|\psi_{\theta}\rangle}{\langle\psi_{\theta}|\psi_{\theta}\rangle},$ (3)
$\mathrm{s.t.}$ $\qquad\psi(\cdots,{x}_{i},\cdots,{x}_{j},\cdots)$
$\qquad=-\psi(\cdots,{x}_{j},\cdots,{x}_{i},\cdots),$
$\qquad\forall i,j\in\{1,2,\cdots,N\},$
and $\theta$ are ansatz parameters for representing neural network wavefunction $\psi_{\theta}$. VMC uses Monte Carlo methods to evaluate the expected value of the total energy
$E_{\mathrm{tot}}$ $=\frac{\langle\psi_{\theta}|\hat{H}|\psi_{\theta}\rangle}{\langle\psi_{\theta}|\psi_{\theta}\rangle}=\int\frac{\hat{H}\psi_{\theta}(x)}{\psi_{\theta}(x)}\frac{\psi_{\theta}^{2}(x)}{\langle\psi_{\theta}|\psi_{\theta}\rangle}\mathop{}\!\mathrm{d}x$ (4)
$=\mathbb{E}_{x\sim p}[E_{L}],$
where $p(x)=\frac{\psi^{2}_{\theta}(x)}{\langle\psi_{\theta}|\psi_{\theta}\rangle}$ and $E_{L}=\frac{\hat{H}\psi_{\theta}(x)}{\psi_{\theta}(x)}$ is so-called local energy. Variational optimization is then performed to find the best approximation of the ground state wavefunction in the ansatz space.
From another perspective, solving the ground state of the time-independent Schrödinger equation can be seen as finding the lowest eigenstate of the Hermitian operator $\hat{H}$, which can be achieved by the power method [72]. By repeatedly applying a propagator to an arbitrary trial state $\psi$, one can eventually get the ground state $\psi_{0}$, as long as $\psi$ and $\psi_{0}$ are not orthogonal. The propagator can be any operator that decays all the other eigenstates of the Hamiltonian $\hat{H}$ while retaining the one with the lowest eigenvalue. When we use the exponential form $e^{-\tau(\hat{H}-E_{T})}$ as propagator, the power method is equivalent to imaginary time evolution under the Wick-rotated Schrödinger equation [73]. Here, $E_{T}<0$ is a scalar close to the ground state energy for normalization.
We consider the linear propagator
$\hat{U}(\tau)=1-\tau(\hat{H}-E_{T}).$ (5)
For the finite time inteval $\tau$, as $n\tau\rightarrow+\infty$, where integer $n$ is the number of time steps, we have $\psi_{0}\propto\hat{U}^{n}(\tau)\psi$.
From the perspective of the power method, VMC can be deemed as an alternating application of two operators: the propagator to evolve towards the ground state, and a projection operator that maps the propagated state back to the ansatz space. Given an ansatz space parametrized by $\theta\in\Omega$, where $\Omega$ is the parameter space, we define the projection operator $\hat{P}_{\mathcal{D}_{1},\mathcal{D}_{2}}({\theta,\eta})$ that projects an arbitrary wavefunction $\phi$ to the ansatz space in the vicinity of $\theta$, with radius $\eta$, while keeping the projected wavefunction normalized, under parameter space metric $\mathcal{D}_{1}$ and Hilbert space metric $\mathcal{D}_{2}$:
$\hat{P}_{\mathcal{D}_{1},\mathcal{D}_{2}}({\theta,\eta})\,\phi=\sqrt{\frac{{\langle\psi_{\theta}|\psi_{\theta}\rangle}}{{\langle\psi_{\theta^{*}}|\psi_{\theta^{*}}\rangle}}}\psi_{\theta^{*}},$ (6)
where
$\theta^{*}=f_{\mathcal{D}_{1},\mathcal{D}_{2}}(\theta,\eta)\phi\triangleq\operatorname*{arg\,min}_{\begin{subarray}{c}\theta^{\prime}\in\Omega\\ \mathcal{D}_{1}(\theta,\theta^{\prime})\leq\eta\end{subarray}}\mathcal{D}_{2}(\phi,\psi_{\theta^{\prime}}).$ (7)
In the case of VMC with natural gradient descent (or stochastic reconfiguration), $\mathcal{D}_{1}$ is the metric induced by the quantum geometric tensor $\mathcal{F}$, and $\mathcal{D}_{2}$ is the Kullback-Leibler (KL) divergence.
At step $n$, the iteration of VMC, starting from parameters $\theta_{n}$, can be written as
${\theta_{n+1}}={\theta_{n}}+\eta_{n}g,$ (8)
$g=\lim_{\tau\rightarrow 0+}\frac{f_{\mathcal{F},\mathrm{KL}}({\theta_{n},\eta_{n}\tau})\,\hat{U}(\tau)\,\psi_{\theta_{n}}-{\theta_{n}}}{\eta_{n}\tau},$
where $\eta_{n}$ is the learning rate, which can vary with the iteration.
Neural network-based variational Monte Carlo [6, 74, 75] utilizes a neural network-based ansatz to parameterize the many-body wavefunction, enabling the accurate capture of complex electronic correlations that are often challenging for traditional quantum chemistry methods. Following Li et al. [9], we use the LapNet ansatz and the Forward Laplacian computational framework.
Instead of directly minimizing total energy, we develop the Lookahead Variational Algorithm (LAVA) for optimization, which is described in more detail below.
4.2 LAVA : the Lookahead Variational Algorithm
LAVA is designed to optimize parameterized wavefunctions by combining the principles of imaginary time evolution (ITE) and variational optimization. We describe LAVA’s implementation in Algorithm 1. Inspired by the Lookahead algorithm [76], our approach calculates the Lookahead direction in Hilbert space following the discretized ITE trajectory (Algorithm 1 line 6-12) and updates the variational ansatz through a projection mechanism ensuring that optimization remains confined to the variational manifold (Algorithm 1 line 13-14).
From this perspective, LAVA “looks ahead” at the space of iteratively propagated states. Instead of directly using the projected propagation for the next step, in iteration $n$, LAVA first constructs temporary state $\psi_{\mathrm{temp}}$ as
$\psi_{\mathrm{temp}}=\hat{P}_{\mathcal{F},\mathrm{KL}}({\theta_{n},\eta_{\mathrm{temp}}})\,\hat{U}(\tau_{\mathrm{temp}})\,\psi_{\theta_{n}},$ (9)
where $\theta_{n}$ denotes the current parameters, and $\eta_{\mathrm{temp}}$, $\tau_{\mathrm{temp}}$ are the projection radius and the small time interval used in this intermediate step, respectively.
From this intermediate $\psi_{\mathrm{temp}}$, LAVA moves along a Lookahead direction:
$\Delta\psi=-\tau\hat{H}(\psi_{\mathrm{temp}}+\psi_{\theta_{n}})+2\tau E_{T}\psi_{\theta_{n}}.$ (10)
Since
$\Delta\psi\propto\frac{1}{2E_{T}}\hat{H}(\psi_{\mathrm{temp}}+\psi_{\theta_{n}})-\psi_{\theta_{n}},$ (11)
to get $\psi_{\theta_{n+1}}$, LAVA applies a projection operator $\hat{P}_{\mathcal{F},\mathrm{SM_{1}}}({\theta_{n},\eta})$ on
$\psi^{\prime}=\frac{1}{E_{T}}\hat{H}(\psi_{\theta_{\mathrm{temp}}}+\psi_{\theta_{n}}).$ (12)
For this projection $\hat{P}_{\mathcal{F},\mathrm{SM_{1}}}({\theta_{n},\eta})$, we use $L_{1}$ score matching (SM):
$\mathrm{SM}_{1}(\phi,\psi;p)=\mathbb{E}_{x\sim p}[\|\nabla_{x}\ln|\phi(x)|-\nabla_{x}\ln|\psi(x)|\|_{1}].$ (13)
Then,
$\psi_{\theta_{n+1}}=\hat{P}_{\mathcal{F},\mathrm{SM_{1}}}({\theta_{n},\eta})\,\psi^{\prime}.$ (14)
For the actual implementation, our projected propagation uses natural gradient descent with Kronecker-Factored Approximate Curvature (known as KFAC) [77], in which a block-diagonal $\mathcal{F}_{\mathrm{KFAC}}$ approximates the quantum geometric tensor $\mathcal{F}$. At iteration $n$, LAVA first performs a VMC step to get intermediate parameters $\theta_{\mathrm{temp}}$:
$\theta_{\mathrm{temp}}=\theta_{n}-\eta_{\mathrm{temp}}\mathcal{F}^{-1}_{\mathrm{KFAC}}g,$ (15)
$g=\mathbb{E}_{x\sim\psi_{\theta_{n}}^{2}}[(E_{L}(x)-E_{\mathrm{tot}})\nabla_{\theta_{n}}\mathrm{ln}|\psi_{\theta_{n}}(x)|],$
where $E_{\mathrm{tot}}$ is the average total energy and $\eta_{\mathrm{temp}}$ is the learning rate for this intermediate step. An unnormalized temporary state $\psi^{\prime}$ is calculated by
$\psi^{\prime}=-\hat{H}\left(\psi_{\theta_{\mathrm{temp}}}\sqrt{\frac{\langle\psi_{\theta_{n}}|\psi_{\theta_{n}}\rangle}{\langle\psi_{\theta_{\mathrm{temp}}}|\psi_{\theta_{\mathrm{temp}}}\rangle}}+\psi_{\theta_{n}}\right).$ (16)
The gradient direction of LAVA then follows
$g_{n}$ $=\nabla_{\theta_{n}}\mathrm{SM}_{1}(\psi^{\prime},\psi_{\theta_{n}};\psi_{\theta_{\mathrm{temp}}}^{2})$ (17)
$=\mathbb{E}_{x\sim\psi_{\theta_{\mathrm{temp}}}^{2}}[\left\langle f(x),\nabla_{\theta_{n}}\nabla_{r}\mathrm{ln}|\psi_{\theta_{n}}(x)|\right\rangle],$
$f(x)=-\mathrm{sgn}\left(\nabla_{r}\ln\left|\frac{\psi^{\prime}(x)}{\psi_{\theta_{n}}(x)}\right|\right).$ (18)
LAVA feeds the gradient into the Adam-KFAC optimizer to update the parameters. Unlike Izadi et al. [78], we directly apply Adam [79] to KFAC preconditioned gradient:
$\theta_{n+1}=\theta_{n}-\mathrm{Adam}(\mathcal{F}_{\mathrm{KFAC}}^{-1}g_{n}).$ (19)
Algorithm 1 LAVA
1:initial parameters $\theta$, samples $\{x^{(i)}\}_{i=1}^{B}$
[…]
5:while $n\leq N$ do
6:  update $x^{(i)}$ by sampling from $\frac{\psi^{2}_{\theta}}{\langle\psi_{\theta}|\psi_{\theta}\rangle}$
7:  $E_{L}(x^{(i)})\leftarrow\frac{\hat{H}\psi_{\theta}(x^{(i)})}{\psi_{\theta}(x^{(i)})}$
8:  $E_{\mathrm{tot}}\leftarrow\frac{1}{B}\sum\limits_{i=1}^{B}E_{L}(x^{(i)})$
9:  $g\leftarrow\frac{1}{B}\sum\limits_{i=1}^{B}(E_{L}(x^{(i)})-E_{\mathrm{tot}})\nabla_{\theta}\mathrm{ln}|\psi_{\theta}(x^{(i)})|$
10:  $\theta_{\mathrm{temp}}\leftarrow\theta-\eta_{\mathrm{temp}}\mathcal{F}^{-1}_{\mathrm{KFAC}}g$
11:  update $x^{(i)}$ by sampling from $\frac{\psi^{2}_{\theta_{\mathrm{temp}}}}{\langle\psi_{\theta_{\mathrm{temp}}}|\psi_{\theta_{\mathrm{temp}}}\rangle}$
12:  $C\leftarrow\sqrt{\frac{\langle\psi_{\theta}|\psi_{\theta}\rangle}{\langle\psi_{\theta_{\mathrm{temp}}}|\psi_{\theta_{\mathrm{temp}}}\rangle}}$
13:  $E_{L}^{\prime}(x^{(i)})\leftarrow\frac{\hat{H}\left(\psi_{\theta}(x^{(i)})+C\psi_{\theta_{\mathrm{temp}}}(x^{(i)})\right)}{\psi_{\theta}(x^{(i)})}$
14:  $g\leftarrow-\frac{1}{B}\sum\limits_{i=1}^{B}\left\langle\mathrm{sgn}(E_{L}^{\prime}(x^{(i)}))\mathrm{sgn}(\nabla_{r^{(i)}}E_{L}^{\prime}(x^{(i)}))\right.,\left.\nabla_{\theta}\nabla_{r^{(i)}}\mathrm{ln}|\psi_{\theta}(x^{(i)})|\right\rangle\end{aligned}$
15:  $\theta\leftarrow\theta-\mathrm{Adam}(\mathcal{F}^{-1}_{\mathrm{KFAC}}g)$
16:  $n\leftarrow n+1$
17:end while
18:return $\theta,\ \{x^{(i)}\}_{i=1}^{B}$
4.3 Neural Scaling Laws
The central premise of our scaling laws establishes that, with LAVA, the energy error of trained models decays following a power law trend with respect to the number of neural network parameters $N_{p}$ governing expressivity:
$E-E_{0}=\alpha N_{p}^{-\beta},$ (20)
where $\alpha$ and $\beta$ are system-dependent variables and $E_{0}$ is the exact ground state energy. We monotonically increase network width and the number of determinants during this scaling-up process. For 111111 systems with reliable experimental benchmarks, the average $r^{2}$ of ordinary least squares (OLS) regression on a logarithmic scale is larger than 0.950.950.95, and residual diagnostics (White test, $p>0.2$) reveals no significant heteroscedasticity. For the details, see Supplementary Note 2.
We also observe the scaling laws of local energy variance $V=\langle(\hat{H}-E)^{2}\rangle$:
$V=\alpha_{v}N_{p}^{-\beta_{v}},$ (21)
where $\alpha_{v}$ and $\beta_{v}$ are system-dependent variables. Under the zero-variance principle of quantum Monte Carlo, our estimated energy approximates the exact ground state energy $E_{0}$ as $V\rightarrow 0$. Since variance estimation requires no exact reference data, unlike error measurement, verifying the relationship is possible for any system where LAVA calculations are available. We confirmed this relationship across various molecular systems by linear regression on a logarithmic scale, and the average $r^{2}$ is larger than 0.970.970.97. Still, OLS is used since the White test shows $p>0.2$. The detailed results are in Supplementary Note 2, and model architectures and training scheme are listed in Supplementary Note 1.1.
4.4 Extrapolation Scheme
For practically trained models, we empirically observed the following relationship between local energy variance and energy:
$E=kV+b,$ (22)
where $k$ and $b$ are system-dependent variables. This relationship is similar to variance-energy extrapolation in Fu et al. [80], which utilizes training data of a fixed model instead of optimized networks of different capacities.
Since theoretically $\mathop{\mathrm{lim}}\limits_{V\rightarrow 0}E=E_{0}$ and empirically $\mathop{\mathrm{lim}}\limits_{N_{p}\rightarrow+\infty}V=0$, we follow Fu et al. [80] and use $b$ as our asymptotic estimates. For various molecular systems, residual diagnostics (White test, $p>0.2$) support the use of ordinary least squares regression, and the average OLS $r^{2}$ for linear regression lines is greater than 0.990.990.99. For systems with experimental reference values, the errors in our extrapolation fall within the experimental uncertainty. Combining neural scaling laws and extrapolation enables the estimation of threshold computational resources to reach sub-kJ/mol accuracy.
[…]
Data Availability
All data supporting the findings of this study are available within the Supplementary Information.
[…]
Acknowledgements
We thank Chenyang Li, Xuefei Xu, Yinan Shu, Donald G. Truhlar for the insight discussion. We thank Zigeng Huang, Qiming Sun, Xiaojie Wu, and the rest of the ByteDance Seed Group for their inspiring ideas and encouragement. We also thank Hang Li for his guidance and support. L.W. and D.H. are supported by National Science and Technology Major Project (2022ZD0114902) and National Science Foundation of China (NSFC92470123, NSFC62276005). W.A.G. thanks the US NSF (CBET 2311117) for support. J.C. is supported by the National Key R&D Program of China (2021YFA1400500) and National Science Foundation of China (12334003).
[…]
Competing interests The authors declare no competing interests.

### v2

@@
 c.
-Cost scaling exponents for N2 as LAVA increases model size and CC methods increase excitation orders.
-For CC, data from Chan et al.
-[24] shows that CC’s computational effort scales as a power-law trend relative to $1/(E_{\mathrm{CC}}-E_{\mathrm{FCI}})$ under cc-pVDZ basis set.
+Cost scaling exponents for N2 as LAVA increases model size and $C$ methods increase excitation orders.
+For $C$, data from Chan et al.
+[24] shows that $C$’s computational effort scales as a power-law trend relative to $1/(E_{\mathrm{CC}}-E_{\mathrm{FCI}})$ under $c$-pVDZ basis set.
 For LAVA, the total GPU time cost scales as a power-law relative to $1/(E_{\mathrm{LAVA}}-E_{\mathrm{SE}})$, where $E_{\mathrm{SE}}$ denotes the LAVA-SE result.
-LAVA maintains a nearly constant exponent along the $N_{2}$ dissociation curve, while CC’s performance deteriorates in the strongly correlated region.
+LAVA maintains a nearly constant exponent along the $N_{2}$ dissociation curve, while $C$’s performance deteriorates in the strongly correlated region.
 d.
@@
 Bars represent the spread of reported values, corresponding to each line, within each method family.
-Stars indicate the improved experimental estimate and the best estimates from systematically improvable theories, namely CCSDTQ/aug-cc-pVTZ in CC, HCI/cc-pVTZ extrapolated to the FCI limit in selected CI, and LAVA in NNQMC.
+Stars indicate the improved experimental estimate and the best estimates from systematically improvable theories, namely CCSDTQ/aug-$c$-pVTZ in $C$, HCI/$c$-pVTZ extrapolated to the FCI limit in selected CI, and LAVA in NNQMC.
 The dashed line shows the value predicted by LAVA, highlighting the consistency among those stars.
@@
 Upper main panel: The ground-state potential energy curve of N2 dissociation.
-The analytic potential functions, namely $\text{MLR}_{3}(9)$, $\text{MLR}_{4}(6,8)$, and $\text{EMO}_{2}(6)$, are fitted from vibrational levels $v=0-19$.
+The analytic potential functions, namely $\text{MLR}_{3}(9)$, $\text{MLR}_{4}(6,8)$, and $\text{EMO}_{2}(6)$, are fitted from vibrational levels v=019v=0-19.
 Upper inset: PECs fitted from different analytic function forms and parameters show a large discrepancy at long bond lengths.
@@
 The previous benchmark, labeled as MLR4(6,8) in Fig.
-3.b, was fitted to experimental vibrational levels $v=0-19$ [39].
+3.b, was fitted to experimental vibrational levels v=019v=0-19 [39].
 These vibrational levels provide a high-accuracy energy benchmark only in the near-equilibrium region, corresponding to bond lengths $r$ between 0.9 to 1.5 $\mathrm{\AA }$ (gray shaded region).
-There are few experimental results for vibrational levels $v$=20-25 [40], albeit with a high uncertainty.
+There are few experimental results for vibrational levels vv=20-25 [40], albeit with a high uncertainty.
 Measuring higher vibrational levels is challenging because the nitrogen dimer becomes unstable in the near-dissociation region.
@@
 At the intermediate region, MLR4(6,8) also shows significant disagreements with multi-reference correlated calculations [41] and previous NNQMC studies [6, 42, 43, 44, 21].
-[…]
+We now provide a definitive benchmark to resolve these discrepancies by performing LAVA calculations across various bond lengths.
+Furthermore, we fit an analytic potential curve in the Morse/long-range form [39], namely MLR3(9), based on LAVA data points which also reproduce the experimental vibrational levels vv=0-19.
+Details on this fitting procedure are given in the Supplementary Note 7.
+This new benchmark MLR3(9) retains the accuracy of the previous benchmark MLR4(6,8) [39] for the near-equilibrium region (0.9Å<r<1.5Å0.9\ $\mathrm{\AA }$<r<1.5\ $\mathrm{\AA }$) and the fully dissociation limit(r>4.0År>4.0\ $\mathrm{\AA }$), while improving the reliability around the near dissociation region.
+In addition, $r_{12}$-MR-ACPF [41, 45], an explicitly correlated multireference method with FCI/CBS accuracy, perfectly aligns with our new benchmark MLR3(9) within $1$ kJ/mol difference, after shifting down by about 5 mHa.
+However, $r_{12}$-MR-ACPF exhibits exponential scaling with respect to the size of reference space.
+This computational complexity necessitates careful optimization of both the reference space and basis set, thereby restricting its practical application to small molecular systems.
+Together, the reproduction of both the available experimental vibrational levels and the relative energies from $r_{12}$-MR-ACPF confirms the accuracy and reliability of our newly fitted benchmark MLR3(9) curve.
 Collectively, the cases in this section demonstrate the reliability and versatility of our approach.
@@
 Basis set convergence of dipole moment of F2O2 for HF and coupled-cluster methods, together with experimental benchmark and LAVA estimates for comparison.
-Dipole moments in Debyes are plotted against correlated-consistent basis sets (cc-pVXZ) and augmented counterparts (aug-cc-pVXZ) for $\mathrm{X}=\mathrm{D},\mathrm{T},\mathrm{Q},5$.
+Dipole moments in Debyes are plotted against correlated-consistent basis sets ($c$-pVXZ) and augmented counterparts (aug-$c$-pVXZ) for $\mathrm{X}=\mathrm{D},\mathrm{T},\mathrm{Q},5$.
 CBS extrapolation scheme follows Supplementary Note 6.2.
@@
 13:  $E_{L}^{\prime}(x^{(i)})\leftarrow\frac{\hat{H}\left(\psi_{\theta}(x^{(i)})+C\psi_{\theta_{\mathrm{temp}}}(x^{(i)})\right)}{\psi_{\theta}(x^{(i)})}$
-14:  $g\leftarrow-\frac{1}{B}\sum\limits_{i=1}^{B}\left\langle\mathrm{sgn}(E_{L}^{\prime}(x^{(i)}))\mathrm{sgn}(\nabla_{r^{(i)}}E_{L}^{\prime}(x^{(i)}))\right.,\left.\nabla_{\theta}\nabla_{r^{(i)}}\mathrm{ln}|\psi_{\theta}(x^{(i)})|\right\rangle\end{aligned}$
+14:  $\begin{aligned} g\leftarrow-\frac{1}{B}\sum\limits_{i=1}^{B}\left\langle\mathrm{sgn}(E_{L}^{\prime}(x^{(i)}))\mathrm{sgn}(\nabla_{r^{(i)}}E_{L}^{\prime}(x^{(i)}))\right.,\left.\nabla_{\theta}\nabla_{r^{(i)}}\mathrm{ln}|\psi_{\theta}(x^{(i)})|\right\rangle\end{aligned}$
 15:  $\theta\leftarrow\theta-\mathrm{Adam}(\mathcal{F}^{-1}_{\mathrm{KFAC}}g)$
@@
 We monotonically increase network width and the number of determinants during this scaling-up process.
-For 111111 systems with reliable experimental benchmarks, the average $r^{2}$ of ordinary least squares (OLS) regression on a logarithmic scale is larger than 0.950.950.95, and residual diagnostics (White test, $p>0.2$) reveals no significant heteroscedasticity.
+For $11$ systems with reliable experimental benchmarks, the average $r^{2}$ of ordinary least squares (OLS) regression on a logarithmic scale is larger than $0.95$, and residual diagnostics (White test, $p>0.2$) reveals no significant heteroscedasticity.
 For the details, see Supplementary Note 2.
@@
 Since variance estimation requires no exact reference data, unlike error measurement, verifying the relationship is possible for any system where LAVA calculations are available.
-We confirmed this relationship across various molecular systems by linear regression on a logarithmic scale, and the average $r^{2}$ is larger than 0.970.970.97.
+We confirmed this relationship across various molecular systems by linear regression on a logarithmic scale, and the average $r^{2}$ is larger than $0.97$.
 Still, OLS is used since the White test shows $p>0.2$.
@@
 [80] and use $b$ as our asymptotic estimates.
-For various molecular systems, residual diagnostics (White test, $p>0.2$) support the use of ordinary least squares regression, and the average OLS $r^{2}$ for linear regression lines is greater than 0.990.990.99.
+For various molecular systems, residual diagnostics (White test, $p>0.2$) support the use of ordinary least squares regression, and the average OLS $r^{2}$ for linear regression lines is greater than $0.99$.
 For systems with experimental reference values, the errors in our extrapolation fall within the experimental uncertainty.
@@
 L.W.
-and D.H.
-are supported by National Science and Technology Major Project (2022ZD0114902) and National Science Foundation of China (NSFC92470123, NSFC62276005).
+is supported by National Science and Technology Major Project (2022ZD0114902) and National Science Foundation of China (NSFC92470123, NSFC62276005).
+D.H.
+is supported by National Science Foundation of China (NSFC62376007).
 W.A.G.